OpenStudy (anonymous):

if l x+y l is smaller than or equal to l x l + l y l prove l x+y+...+xn l is smaller than or equal to l x l + l y l + ... + l xn l where n is Natural number

OpenStudy (anonymous):

the "if" part probably means you are supposed to do this by induction right?

OpenStudy (anonymous):

yes by PIM

OpenStudy (anonymous):

whatever that means in any case do you know how to do a proof by induction?

OpenStudy (anonymous):

umm not quite

OpenStudy (anonymous):

I know how to do it when there is equal sign and without absolute value

OpenStudy (anonymous):

ok first step is base case ,but you are given that as an "if " part, ie. you get to assume \[|x_1+x_2|\leq |x_1|+|x_2|\] normally you would have to prove that first

OpenStudy (anonymous):

squaring each side?

OpenStudy (anonymous):

there are lots of proofs, that is one of them, but you don't need to do it that came under the heading of IF normally you would, but you are spared that here because of the way the question was written

OpenStudy (anonymous):

oh ok thx then how would you proof other parts? I mean the text book itself says to set that x is S and prove them by x+1

OpenStudy (anonymous):

then you get to ASSUME it is true if \(n=k\) i.e you assume \[|x_1+x_2+...+x_k|\leq |x_1|+|x_2|+...+|x_k|\] and now show it is true for \(n=k+1\)

OpenStudy (anonymous):

if we put \(n=k+1\) we are trying to show that \[|x_1+x_2+...+x_k+x_{k+1}|\leq |x_1|+|x_2|+...+|x_k|+|x_{k+1}|\]

OpenStudy (anonymous):

break part in to two pieces so you can use \[|x_1+x_2+...+x_k|\leq |x_1|+|x_2|+...+|x_k|\] just maybe put parentheses

OpenStudy (anonymous):

put parentheses on where?

OpenStudy (anonymous):

\[|(x_1+x_2+...+x_k)+x_{k+1}|\]

OpenStudy (anonymous):

you get to assume (because after all \((x_1+x_2+...+x_k)\) is just some number) that \[|(x_1+x_2+...+x_k)+x_{k+1}|\leq |(x_1+x_2+...+x_k)|+|x_{k+1}|\]

OpenStudy (anonymous):

then the "induction hypothesis" will finish it, because you get to assume that \[|x_1+x_2+...+x_k|\leq |x_1|+|x_2|+...+|x_k|\]

OpenStudy (anonymous):

thank you it helped me alot

OpenStudy (anonymous):

but do you need closure for this?

OpenStudy (anonymous):

yw with these problems it is best to write exactly what the induction hypothesis is, so you can figure out how to get back to it

OpenStudy (anonymous):

closure? lost me there

OpenStudy (anonymous):

Like assuming that (x 1 +x 2 +...+x k ) to some number in the end of the proof

OpenStudy (anonymous):

no you don't have to worry about it being a number it is a number